# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v1

@@ 6 Conclusion
This paper advances the field of hyperbolic LVMs by augmenting the hyperbolic manifold with a Riemannian pullback metric that combines the hyperbolic geometry with the geometry of the data distribution. By minimizing the curve energy on the pullback metric, we computed geodesics that adhere to the data manifold in the hyperbolic latent space. To do so, we addressed the limitations of current auto-differentiation techniques by providing analytical solutions for the hyperbolic heat kernel derivatives. Via multiple experiments, we demonstrated that pullback geodesics outperform standard hyperbolic geodesics.
It is worth noting that the benefits of pullback geodesics are most evident when the data exhibits smooth transitions. They become less effective when the data is inherently comprised of distinct clusters, as the pullback geodesics cannot cross high-energy regions, i.e., the data manifold boundaries. Future work could focus on generating latent embeddings with smoother transitions. Additionally, the need for computing manual kernel derivatives is unsatisfactory. Future work will explore alternative autodifferentiation techniques built on KeOps (Feydy et al.,, 2020) to overcome this practical issue. Finally, the approximation of the 2D hyperbolic heat kernel is computationally expensive. Performance could potentially be increased by exploring different sampling strategies, e.g., by sampling from a Rayleigh distribution rather than from a Gaussian.
@@ Acknowledgements
This work was supported by the Carl Zeiss Foundation under the project JuBot and the European Union’s Horizon Europe Framework Programme under grant agreement No 101070596 (euROBIN).
@@ Appendix A Hyperbolic Manifold
where $\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}=\bm{x}^{\mathsf{T}}\bm{G}\bm{y}$ is the lorentzian inner product with the Lorentzian metric tensor $\bm{G}=\text{diag}(-1,1,...,1)$. Note that, since the curvature of the hyperbolic manifold is constant, the metric tensor does not depend on $\bm{x}$ as for general manifolds, i.e., $\bm{G}_{\bm{x}}=\bm{G}$ for all $\bm{x}\in\mathbb{H}^{D}_{\mathcal{L}}$. For the hyperbolic manifold, closed form solutions exist for the standard manifold operations. Specifically, the geodesic distance, exponential map, logarithmic map, parallel transport, and tangent space projection are given as,
@@
$\mathrm{proj}_{\bm{x}}(\bm{w})$ $=\bm{P}_{\bm{x}}\bm{g}=(\bm{G}+\bm{x}\bm{x}^{\mathsf{T}})\bm{w}\,.$ (35)
These operations are illustrated by Figure 5.
@@ Appendix B Hyperbolic Kernel Derivatives
This section provides the complete derivations of the hyperbolic heat kernel derivatives that are necessary to compute the pullback metric tensor and to optimize geodesics on it. Equation (20) defines the pullback metric tensor $\bm{G}_{\bm{x}^{*}}$ based on the Jacobian mean $\bm{\mu}_{J}$ (17) and covariance matrix $\bm{\Sigma}_{J}$ (18), which in turn require the kernel derivatives $\partial k(\bm{x}^{*},\bm{X})=\frac{\partial}{\partial\bm{x}^{*}}k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N}$ and $\partial^{2}k(\bm{x}^{*},\bm{x}^{*})=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}\in\mathbb{R}^{D_{x}\times D_{x}}$. Additionally, minimizing the curve energy (21), requires the derivative of the pullback metric tensor, i.e., the derivative of $\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J}$, which requires the kernel derivative $\dfrac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})$ and $\dfrac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})$. Specifically, we have
$\dfrac{\partial}{\partial\bm{x}^{*}}(\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J})=\left[\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}_{J}^{\mathsf{T}}\right]\times_{1}\bm{\mu}_{J}+\left[\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}_{J}\right]\times_{0}\bm{\mu}_{J}+D_{y}\left[\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}\right],$ (37)
with $\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}_{J}^{\mathsf{T}}=\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\right]\times_{1}(\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N})^{-1}\bm{Y},$ (38)
$\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}=\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial^{2}k(\bm{x}^{*},\bm{x}^{*})\right]-\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\right]\times_{1}\bm{S}_{J}^{\mathsf{T}}-\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{X},\bm{x}^{*})\right]\times_{0}\bm{S}_{J}^{\mathsf{T}},$ (39)
@@
where we have $\bm{\mu}_{J}\in\mathbb{R}^{D_{y}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}\in\mathbb{R}^{D_{y}\times D_{x}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N\times D_{x}}$, and ×$\times_{n}$ denotes the $n$-th mode tensor product (Kolda and Bader,, 2009). Next, we provide the details of the derivation of the kernel derivatives for $\mathbb{H}^{2}_{\mathcal{L}}$ and $\mathbb{H}^{3}_{\mathcal{L}}$.
B.1 2D Hyperbolic Kernel Derivatives
As discussed in Section 4.1, the derivatives of the 2D hyperbolic heat kernel (6) are entirely determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$. Equations (23), (24), (25), and (26) show that these derivatives in turn are defined by the first and second derivatives of $\phi_{l}(\bm{x}_{\mathcal{P}})=e^{(1+2s_{l}i)\langle\bm{x}_{\mathcal{P}},\bm{b}_{l}\rangle}$, which are computed as
@@
$\qquad\qquad\qquad+\phi_{l}(\bm{x}_{\mathcal{P}})\left[\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\bm{x}_{\mathcal{P}}\right]\times_{0}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right]$
@@
$\qquad\qquad\qquad+\left[\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\bm{x}_{\mathcal{P}}\right]\times_{0}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right]$
@@
B.2 3D Hyperbolic Kernel Derivatives
This section provides the derivations of the 3D hyperbolic heat kernel derivatives. Section 4.2 introduced the helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$ and $s=\sqrt{u^{2}-1}$ with derivatives $\frac{\partial u}{\partial\bm{x}}=\bm{G}\bm{z}$, $\frac{\text{d}\rho}{\text{d}u}=\frac{-1}{s}$, and $\frac{\text{d}s}{\text{d}u}=\frac{u}{s}$. Using these expressions, the first and second derivatives of the 3D hyperbolic heat kernel $k(\bm{x},\bm{y})=\frac{\rho}{s}\,e^{-\frac{\rho^{2}}{\kappa}}$ are given as,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}\rho\right]\dfrac{1}{s}e^{-\frac{\rho^{2}}{\kappa}}+\rho\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}s^{-1}\right]e^{-\frac{\rho^{2}}{\kappa}}+\dfrac{\rho}{s}\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\kappa}}\right],$ (47)
$=-\dfrac{1}{s^{2}}\bm{G}\bm{z}e^{-\frac{\rho^{2}}{\kappa}}-\dfrac{\rho u}{s^{3}}\bm{G}\bm{z}e^{-\frac{\rho^{2}}{\kappa}}+\dfrac{\rho}{s^{2}}\bm{G}\bm{z}\dfrac{2\rho}{\kappa}e^{-\frac{\rho^{2}}{\kappa}},$ (48)
$=g(u)e^{-\frac{\rho^{2}}{\kappa}}\bm{G}\bm{z}\,,$ (49)
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\bm{Gz}\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\kappa}}+\bm{Gz}g(u)\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\kappa}}\right]+g(u)e^{-\frac{\rho^{2}}{\kappa}}\left[\dfrac{\partial}{\partial\bm{z}}\bm{Gz}\right],$ (50)
$=\left[\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\kappa}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}+g(u)\dfrac{-1}{s}\dfrac{-2\rho}{\kappa}e^{-\frac{\rho^{2}}{\kappa}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}+g(u)e^{-\frac{\rho^{2}}{\kappa}}\bm{G},$ (51)
$=\left(h(u)\bm{G}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}+g(u)\bm{G}\right)e^{-\frac{\rho^{2}}{\kappa}}\,,$ (52)
$\dfrac{\partial^{2}}{\partial\bm{x}\partial\bm{x}}k(\bm{x},\bm{z})$ $=h(u)\bm{G}\bm{z}\bm{z}^{\mathsf{T}}\bm{G}e^{-\frac{\rho^{2}}{\kappa}}\,.$ (53)
@@
$\dfrac{\partial}{\partial\bm{z}}\bm{K}_{j}$ $=\dfrac{\partial}{\partial\bm{z}}\big{(}h(u)\bm{G}\bm{z}x_{j}+g(u)\bm{e}_{j}\big{)}\Delta_{1j}e^{-\frac{\rho^{2}}{\kappa}},$ (54)
@@
$=\left(\left[\dfrac{\text{d}}{\text{d}u}h(u)\right]\bm{G}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}x_{j}+h(u)\bm{G}x_{j}+\left[\dfrac{\text{d}}{\text{d}u}g(u)\right]\bm{e}_{j}\bm{x}^{\mathsf{T}}\bm{G}\right)\Delta_{1j}e^{-\frac{\rho^{2}}{\kappa}}$ (55)
$\quad+\big{(}h(u)\bm{G}\bm{z}x_{j}+g(u)\bm{e}_{j}\big{)}\Delta_{1l}\dfrac{2\rho}{\kappa s}e^{-\frac{\rho^{2}}{\kappa}}\bm{x}^{\mathsf{T}}\bm{G},$ (56)
$=\left(q(u)\bm{G}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}x_{j}+h(u)\bm{G}x_{j}+h(u)\bm{e}_{j}\bm{x}^{\mathsf{T}}\bm{G}\right)\Delta_{1j}e^{-\frac{\rho^{2}}{\kappa}}\,,$ (57)
@@
$\dfrac{\partial}{\partial\bm{x}}\bm{K}_{i}$ $=\left(q(u)\bm{G}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}z_{i}\,+h(u)\bm{G}z_{i}\,+h(u)\bm{Gz}\bm{e}_{i}^{\mathsf{T}}\right)\Delta_{1i}e^{-\frac{\rho^{2}}{\kappa}}\,.$ (58)

### v2

@@ 6 Conclusion
This paper advances the field of hyperbolic LVMs by augmenting the hyperbolic manifold with a Riemannian pullback metric that combines the hyperbolic geometry with the geometry of the data distribution. By minimizing the curve energy on the pullback metric, we computed geodesics that adhere to the data manifold in the hyperbolic latent space. To do so, we addressed the limitations of current auto-differentiation techniques by providing analytical solutions for the hyperbolic SE kernel derivatives. Via multiple experiments, we demonstrated that pullback geodesics outperform standard hyperbolic geodesics.
It is worth noting that the benefits of pullback geodesics are most evident when the data exhibits smooth transitions. They become less effective when the data is inherently comprised of distinct clusters, as the pullback geodesics cannot cross high-energy regions, i.e., the data manifold boundaries. This paper focused on 2D and 3D hyperbolic latent spaces as they are easier to visualize and computationally less expensive than higher-dimensional latent spaces, while benefiting from the increased capacity of hyperbolic spaces to embed hierarchical data. Future work will investigate extensions to higher-dimensional hyperbolic latent spaces, which require to deal with more complex expressions for hyperbolic kernels. Additionally, the need for computing manual kernel derivatives is unsatisfactory. Future work will explore alternative autodifferentiation techniques built on KeOps (Feydy et al., 2020) to overcome this practical issue. Finally, the approximation of the 2D hyperbolic SE kernel is computationally expensive. Performance could potentially be increased by exploring different sampling strategies, e.g., by sampling from a Rayleigh distribution rather than from a Gaussian.
@@ Acknowledgements
This work was partially supported by the Carl Zeiss Foundation under the project JuBot, by the European Union’s Horizon Europe Framework Programme under grant agreement No 101070596 (euROBIN), and by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ Appendix A Hyperbolic Manifold
where $\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}=\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}\bm{y}$ is the Lorentzian inner product with the Lorentzian metric tensor $\bm{G}^{\mathcal{L}}=\text{diag}(-1,1,...,1)$. Note that, since the curvature of the hyperbolic manifold is constant, the metric tensor does not depend on $\bm{x}$ as for general manifolds, i.e., $\bm{G}^{\mathcal{L}}_{\bm{x}}=\bm{G}^{\mathcal{L}}$ for all $\bm{x}\in\mathbb{H}^{D}_{\mathcal{L}}$. For the hyperbolic manifold, closed form solutions exist for the standard manifold operations. Specifically, the geodesic distance, exponential map, logarithmic map, parallel transport, and tangent space projection are given as,
@@
$\mathrm{proj}_{\bm{x}}(\bm{w})$ $=\bm{P}_{\bm{x}}\bm{w}=(\bm{G}^{\mathcal{L}}+\bm{x}\bm{x}^{\mathsf{T}})\bm{w}\,.$ (37)
These operations are illustrated by Fig. 6.
@@
As mentioned in the main text, we mostly rely on the Lorentz model $\mathbb{H}^{D}_{\mathcal{L}}$ in our algorithms, but use the Poincaré model for visualization, as it provides a more intuitive representation. The Poincaré model is defined as
$\mathbb{H}^{D}_{\mathcal{P}}=\{\bm{x}_{\mathcal{P}}\in\mathbb{R}^{D}\>|\>\|\bm{x}_{\mathcal{P}}\|<1\}.$ (39)
It is possible to map points from the Lorentz to the Poincaré model via an isometric mapping $f:\mathbb{H}^{D}_{\mathcal{L}}\to\mathbb{H}^{D}_{\mathcal{P}}$ such that
$f(\bm{x})=\frac{\left(x_{1},\dots,x_{d}\right)^{\mathsf{T}}}{x_{0}+1},$ (40)
where $\bm{x}\in\mathbb{H}^{D}_{\mathcal{L}}$ with components $x_{0},x_{1},\dots,x_{d}$. The inverse mapping $f^{-1}:\mathbb{H}^{D}_{\mathcal{P}}\to\mathbb{H}^{D}_{\mathcal{L}}$ is defined as follows
$f^{-1}(\bm{y})=\frac{\left(1+\|\bm{y}\|^{2},2y_{1},\dots,2y_{d}\right)^{\mathsf{T}}}{1-\|\bm{y}^{2}\|},$ (41)
with $\bm{y}\in\mathbb{H}^{D}_{\mathcal{P}}$ with components $y_{1},\dots,y_{d}$.
@@ Appendix B Hyperbolic Kernel Derivatives
This section provides the complete derivations of the hyperbolic SE kernel derivatives that are necessary to compute the pullback metric tensor and to optimize geodesics on it. Eq. (22) defines the pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ based on the Jacobian mean $\bm{\mu}_{J}$ (19) and covariance matrix $\bm{\Sigma}_{J}$ (20), which in turn require the kernel derivatives $\partial k(\bm{x}^{*},\bm{X})=\frac{\partial}{\partial\bm{x}^{*}}k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N}$ and $\partial^{2}k(\bm{x}^{*},\bm{x}^{*})=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}\in\mathbb{R}^{D_{x}\times D_{x}}$. Additionally, minimizing the curve energy (23), requires the derivative of the pullback metric tensor, i.e., the derivative of $\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J}$, which requires the kernel derivative $\dfrac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})$ and $\dfrac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})$. Specifically, we have
$\dfrac{\partial}{\partial\bm{x}^{*}}(\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J})=\left[\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}_{J}^{\mathsf{T}}\right]\times_{2}\bm{\mu}_{J}^{\mathsf{T}}+\left[\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}_{J}\right]\times_{1}\bm{\mu}_{J}^{\mathsf{T}}+D_{y}\left[\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}\right],$ (42)
with $\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}_{J}^{\mathsf{T}}=\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\right]\times_{2}\bm{Y}^{\mathsf{T}}(\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N})^{-1},$ (43)
$\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}=\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial^{2}k(\bm{x}^{*},\bm{x}^{*})\right]-\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\right]\times_{2}\bm{S}_{J}-\left[\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{X},\bm{x}^{*})\right]\times_{1}\bm{S}_{J},$ (44)
@@
where we have $\bm{\mu}_{J}\in\mathbb{R}^{D_{y}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\bm{\mu}\in\mathbb{R}^{D_{y}\times D_{x}\times D_{x}}$, $\dfrac{\partial}{\partial\bm{x}^{*}}\partial k(\bm{x}^{*},\bm{X})\in\mathbb{R}^{D_{x}\times N\times D_{x}}$, and ×$\times_{n}$ denotes the $n$-th mode tensor product (Kolda & Bader, 2009a). Next, we provide the details of the derivation of the kernel derivatives for $\mathbb{H}^{2}_{\mathcal{L}}$ and $\mathbb{H}^{3}_{\mathcal{L}}$.
B.1 2D Hyperbolic SE Kernel Derivatives
As discussed in Sec. 4.1, the derivatives of the 2D hyperbolic SE kernel (6) are entirely determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$. Eqs (25), (26), (27), and (28) show that these derivatives in turn are defined by the first and second derivatives of $\phi_{l}(\bm{x}_{\mathcal{P}})=e^{(1+2s_{l}i)\langle\bm{x}_{\mathcal{P}},\bm{b}_{l}\rangle}$, which are computed as
@@
$\qquad\qquad\qquad+\phi_{l}(\bm{x}_{\mathcal{P}})\left[\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\bm{x}_{\mathcal{P}}\right]\times_{1}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right]$
@@
$\qquad\qquad\qquad+\left[\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\bm{x}_{\mathcal{P}}\right]\times_{1}\left[\dfrac{\partial}{\partial\bm{x}_{\mathcal{P}}}\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\right]$
@@
B.2 3D Hyperbolic SE Kernel Derivatives
This section provides the derivations of the 3D hyperbolic SE kernel derivatives. Sec. 4.2 introduced the helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$ and $s=\sqrt{u^{2}-1}$ with derivatives $\frac{\partial u}{\partial\bm{x}}=\bm{G}^{\mathcal{L}}\bm{z}$, $\frac{\text{d}\rho}{\text{d}u}=\frac{-1}{s}$, and $\frac{\text{d}s}{\text{d}u}=\frac{u}{s}$. Using these expressions, the first and second derivatives of the 3D hyperbolic SE kernel $k(\bm{x},\bm{y})=\frac{\rho}{s}\,e^{-\frac{\rho^{2}}{\nu}}$ are given as,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}\rho\right]\dfrac{1}{s}e^{-\frac{\rho^{2}}{\nu}}+\rho\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}}{\text{d}u}s^{-1}\right]e^{-\frac{\rho^{2}}{\nu}}+\dfrac{\rho}{s}\left[\dfrac{\partial u}{\partial\bm{x}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\nu}}\right],$ (52)
$=-\dfrac{1}{s^{2}}\bm{G}^{\mathcal{L}}\bm{z}e^{-\frac{\rho^{2}}{\nu}}-\dfrac{\rho u}{s^{3}}\bm{G}^{\mathcal{L}}\bm{z}e^{-\frac{\rho^{2}}{\nu}}+\dfrac{\rho}{s^{2}}\bm{G}^{\mathcal{L}}\bm{z}\dfrac{2\rho}{\nu}e^{-\frac{\rho^{2}}{\nu}},$ (53)
$=g(u)e^{-\frac{\rho^{2}}{\nu}}\bm{G}^{\mathcal{L}}\bm{z}\,,$ (54)
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\bm{Gz}\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\nu}}+\bm{Gz}g(u)\left[\dfrac{\partial u}{\partial\bm{z}}\dfrac{\text{d}\rho}{\text{d}u}\dfrac{\text{d}}{\text{d}\rho}e^{-\frac{\rho^{2}}{\nu}}\right]+g(u)e^{-\frac{\rho^{2}}{\nu}}\left[\dfrac{\partial}{\partial\bm{z}}\bm{Gz}\right],$ (55)
$=\left[\dfrac{\text{d}}{\text{d}u}g(u)\right]e^{-\frac{\rho^{2}}{\nu}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)\dfrac{-1}{s}\dfrac{-2\rho}{\nu}e^{-\frac{\rho^{2}}{\nu}}\bm{Gz}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)e^{-\frac{\rho^{2}}{\nu}}\bm{G}^{\mathcal{L}},$ (56)
$=\left(h(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)\bm{G}^{\mathcal{L}}\right)e^{-\frac{\rho^{2}}{\nu}}\,,$ (57)
$\dfrac{\partial^{2}}{\partial\bm{x}\partial\bm{x}}k(\bm{x},\bm{z})$ $=h(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{z}^{\mathsf{T}}\bm{G}^{\mathcal{L}}e^{-\frac{\rho^{2}}{\nu}}\,.$ (58)
@@
$\dfrac{\partial}{\partial\bm{z}}\bm{K}_{j}$ $=\dfrac{\partial}{\partial\bm{z}}\big{(}h(u)\bm{G}^{\mathcal{L}}\bm{z}x_{j}+g(u)\bm{e}_{j}\big{)}\Delta_{1j}e^{-\frac{\rho^{2}}{\nu}},$ (59)
@@
$=\left(\left[\dfrac{\text{d}}{\text{d}u}h(u)\right]\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}x_{j}+h(u)\bm{G}^{\mathcal{L}}x_{j}+\left[\dfrac{\text{d}}{\text{d}u}g(u)\right]\bm{e}_{j}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}\right)\Delta_{1j}e^{-\frac{\rho^{2}}{\nu}}$ (60)
$\quad+\big{(}h(u)\bm{G}^{\mathcal{L}}\bm{z}x_{j}+g(u)\bm{e}_{j}\big{)}\Delta_{1l}\dfrac{2\rho}{\nu s}e^{-\frac{\rho^{2}}{\nu}}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}},$ (61)
$=\left(q(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}x_{j}+h(u)\bm{G}^{\mathcal{L}}x_{j}+h(u)\bm{e}_{j}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}\right)\Delta_{1j}e^{-\frac{\rho^{2}}{\nu}}\,,$ (62)
@@
$\dfrac{\partial}{\partial\bm{x}}\bm{K}_{i}$ $=\left(q(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}z_{i}\,+h(u)\bm{G}^{\mathcal{L}}z_{i}\,+h(u)\bm{Gz}\bm{e}_{i}^{\mathsf{T}}\right)\Delta_{1i}e^{-\frac{\rho^{2}}{\nu}}\,.$ (63)